OpenStudy (anonymous):

The improper fraction 38/9 is equivalent to the mixed number 4 2/9. A) True B) False

OpenStudy (whpalmer4):

If you have an mixed number A + B/C (in this case, A = 4, B =2, C = 9) that is equivalent to the improper fraction (A*C + B)/C

OpenStudy (whpalmer4):

given that, do you think it is true, or false?

OpenStudy (anonymous):

True I think

OpenStudy (whpalmer4):

that should be "True! Definitely!" :-) 4*9 = 36 36+2 = 38

OpenStudy (anonymous):

Haha thank you! For helping me!

OpenStudy (whpalmer4):

remember, the whole number gets multiplied by 1 (which doesn't change its value) but that is 1 written as denominator/denominator: 4 * 9/9 = 36/9 then we add the fraction: 2/9 36/9 + 2/9 = (36+2)/9 = 38/9 Easy peasy :-)

OpenStudy (whpalmer4):

you're welcome!

OpenStudy (anonymous):

that's trueand so the answer is A which is true

OpenStudy (angelab97):

true